OpenStudy (anonymous):

Suppose that the probability of a person having a headache is 0.01, that the probability of a person having a fever given that the person has a headache is 0.4, and that the probability of a person having a fever is 0.02. find the probability that a person has a headache given that the person has a fever

OpenStudy (mathmate):

 \( P(H)=0.01 \) \( P(F)=0.02 \) \( P(F|H)=0.4 = P(F \cap H)/P(H) \) => \( P(F \cap H)=0.4P(H)=0.004\) \( P(H|F) = P(F\cap H)/P(H) = 0.004/0.02 = 0.2 \)